# Inference-Time Alignment for Diffusion Models via Doob’s Matching

### v1

@@ 3.4  A summary of procedure
where $0\leq k\leq K-1$, the pre-trained base score estimator $\widehat{\mathbf{s}}$ is defined as (2.5), and the Doob’s guidance estimator $\widehat{\mathbf{g}}^{\lambda}$ is defined as (3.22).
We apply post-processing to the generated particle $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$ to ensure numerical stability and facilitate the theoretical analysis presented in Theorem 4.5 (Lee et al., 2023a; Chen et al., 2023). First, we assume the target distribution $q_{0}$ is concentrated on a domain centered at the origin, such as a distribution with compact support (Assumption 1) or with light tails. Consequently, we introduce a truncation operator to the particles obtained from the controllable diffusion model (3.23). Second, because the controllable diffusion process is terminated at an early-stopping time $T_{0}$, there exists a mean shift between the target distribution $q_{0}$ and the early-stopping distribution $q_{T_{0}}\approx\widehat{q}_{T-T_{0}}$, as indicated by (2.2). To mitigate this drift, we employ a scaling operator. Specifically, for $R>0$, we define a truncation operator ${\mathcal{T}}_{R}:\mathbf{z}\mapsto\mathbf{z}{\mathbbm{1}}_{B({\mathbf{0}},R)}(\mathbf{z})$ and a scaling operator ${\mathcal{M}}:\mathbf{z}\mapsto\mu_{T_{0}}^{-1}\mathbf{z}$. The final processed particle is defined as
@@
Input: Base score estimator $\widehat{\mathbf{s}}$, the weight function $w$, the regularization parameter $\lambda$, the step size $h$, and the number of steps $K$.
@@
2 Estimate Doob’s $h$-function by $\widehat{h}_{t}^{\lambda}$ via gradient penalized Doob’s matching (3.21).
@@
6   Evaluate the base score: $\widehat{\mathbf{s}}_{k}\leftarrow\widehat{\mathbf{s}}(kh,\widehat{\mathbf{Z}}_{kh}^{\leftarrow})$.
@@
Algorithm 1 Inference-time alignment via Doob’s matching
@@ 4  Convergence Analysis
In this section, we establish a non-asymptotic convergence rate analysis for Doob’s matching (3.21) and the induced controllable diffusion model (3.23) introduced in Section 3.
@@ 4.1  Assumptions
The support of the target distribution $q_{0}$ is a compact set contained within the hypercube $\{\mathbf{x}\in{\mathbb{R}}^{d}:\|\mathbf{x}\|_{\infty}\leq 1\}$.
@@
Assumption 2 implies that the base distribution $p_{0}$ and the tilted distribution $q_{0}$ satisfy mutual absolute continuity, ensuring that their supports coincide (i.e., $\mathrm{supp}(p_{0})=\mathrm{supp}(q_{0})$). This condition is crucial for establishing the regularity of Doob’s $h$-function. Furthermore, the ratio $\kappa\coloneq\bar{B}/\underline{B}$ serves as a condition number that characterizes the difficulty of the controllable diffusion task, as discussed in the context of posterior sampling by Purohit et al. (2025); Ding et al. (2024b); Chang et al. (2025b).
Assumption 3 (Base score estimation error).
The base score estimator $\widehat{\mathbf{s}}$ defined in (2.5) satisfies the following error bound:
$\frac{1}{T}\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{P}}}\left[\|\widehat{\mathbf{s}}(kh,\mathbf{X}_{kh}^{\leftarrow})-\nabla\log p_{T-kh}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right]\leq\varepsilon_{\mathrm{base}}^{2}.$
Assumption 3 requires the $L^{2}$-error of the base score estimator $\widehat{\mathbf{s}}$ to be bounded with respect to the base path measure $\mathbb{P}$}. In our setting, where numerous samples from the base distribution $p_{0}$ are available, estimators satisfying this bound can be obtained via implicit score matching (Hyvärinen, 2005), sliced score matching (Song et al., 2020), or denoising score matching (Vincent, 2011). While one can derive explicit bounds of base score matching as Tang and Yang (2024); Oko et al. (2023); Fu et al. (2024); Ding et al. (2024a); Yakovlev and Puchkin (2025a); Yakovlev et al. (2025) using non-parametric regression theory for deep neural networks (Bauer and Kohler, 2019; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Jiao et al., 2023), we adopt this condition to maintain clarity of presentation, following the convention of Lee et al. (2023a); Chen et al. (2023); Beyler and Bach (2025); Kremling et al. (2025).
4.2  Regularity of Doob’s $h$-function
In this subsection, we establish the regularity properties of Doob’s $h$-function defined in (3.10).
@@ 4.2  Regularity of Doob’s $h$-function
The proof of Proposition 4.1 is deferred to Appendix C.1. It is worth noting that Proposition 4.1 relies solely on the boundedness of the weight function $w$, without requiring the existence or smoothness of its gradients. Nevertheless, we establish that Doob’s $h$-function admits bounded derivatives. This result stems from the definition of Doob’s $h$-function as a posterior expectation under a Gaussian likelihood; the inherent smoothness of the Gaussian kernel endows the posterior expectation with strong regularity properties.
@@ 4.3  Error bounds for the Doob’s guidance estimator
We begin by introducing the concept of Rademacher complexity (Bartlett and Mendelson, 2002; Mohri et al., 2018), which is crucial for analyzing the generalization error.
Definition 1 (Rademacher complexity).
$\widehat{{\mathfrak{R}}}({\mathscr{H}}|\mathbf{X}^{1:n})\coloneq{\mathbb{E}}\Big[\sup_{h\in{\mathscr{H}}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h(\mathbf{X}^{i})\Big|\mathbf{X}^{1:n}\Big],$
For notational simplicity, we define the derivative classes and their associated Rademacher complexity as follows:
(4.1) ${\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t})\coloneq\sum_{k=1}^{d}{\mathfrak{R}}_{n}(D_{k}{\mathscr{H}}_{t}),\quad D_{k}{\mathscr{H}}_{t}\coloneq\big\{D_{k}h_{t}:h_{t}\in{\mathscr{H}}_{t}\big\}.$
The following lemma provides an oracle inequality for the Doob’s matching.
@@
Suppose Assumptions 1 and 2 hold. Let $t\in(0,T)$ and let ${\mathscr{H}}_{t}$ be a convex hypothesis class. Let $\widehat{h}_{t}^{\lambda}$ be the gradient penalized empirical risk minimizer defined as (3.21), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequalities hold:
$\lesssim\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{H^{1}(p_{T-t})}^{2}}_{\text{approximation}}+\underbrace{\bar{B}{\mathfrak{R}}_{n}({\mathscr{H}}_{t})+\frac{\lambda\bar{B}}{\sigma_{T-t}^{2}}{\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t})}_{\text{generalization}}+\underbrace{\frac{\lambda^{2}d^{2}\bar{B}^{2}}{\sigma_{T-t}^{8}}}_{\text{regularization}},$
$\lesssim\underbrace{\frac{1}{\lambda}\inf_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{H^{1}(p_{T-t})}^{2}}_{\text{approximation}}+\underbrace{\frac{\bar{B}}{\lambda}{\mathfrak{R}}_{n}({\mathscr{H}}_{t})+\frac{\bar{B}}{\sigma_{T-t}^{2}}{\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t})}_{\text{generalization}}+\underbrace{\frac{\lambda d^{2}\bar{B}^{2}}{\sigma_{T-t}^{8}}}_{\text{regularization}},$
@@
The proof of Lemma 4.2 is deferred to Appendix C.2. Both oracle inequalities for Doob’s $h$-function and its gradient decompose the error into three components: approximation error, generalization error, and regularization error.
@@
    The approximation error is defined as the minimal $H^{1}$-distance between functions in the hypothesis class ${\mathscr{H}}_{t}$ and the target function $h_{t}^{*}$, measuring the approximation capability of ${\mathscr{H}}_{t}$.
@@
    The regularization error is introduced by the gradient penalty in the objective functional, which causes the minimizer of the population risk to deviate from the exact Doob’s $h$-function $h_{t}^{*}$ (3.10).
@@
Lemma 4.2 is analogous to the oracle inequality found in regression problems. Let $\widehat{h}_{t}$ be the estimator of the Doob’s $h$-function obtained via vanilla regression (3.18). Informally, the following oracle inequality holds:
(4.2) ${\mathbb{E}}\Big[\|\widehat{h}_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\lesssim\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}}_{\text{approximation}}+\underbrace{\bar{B}{\mathfrak{R}}_{n}({\mathscr{H}}_{t})}_{\text{generalization}}.$
@@
    In vanilla regression, the generalization error in (4.2) depends only on the complexity of the hypothesis class. In contrast, the generalization error bounds in Lemma 4.2 also depend on the complexity of the derivative classes (4.1). This occurs because the objective functional of the gradient penalized Doob’s matching (3.19) includes the gradient norm term. Consequently, the error from finite-sample approximation is influenced not only by the complexity of the hypothesis class but also by that of the derivative class (4.1).
@@
    The most significant difference lies in the regularization error. If we focus solely on the oracle inequality for $\widehat{h}_{t}^{\lambda}$, letting $\lambda$ go to zero reduces the expression to the vanilla regression oracle inequality (4.2). However, the bound for the gradient $\nabla\widehat{h}_{t}^{\lambda}$ diverges as the regularization parameter $\lambda$ approaches zero. This highlights the key advantage of our gradient penalized method: the gradient penalty is essential for guaranteeing simultaneous convergence of both the estimator value and its gradient. Additionally, there exists a trade-off with respect to $\lambda$ in the oracle inequality for $\widehat{h}_{t}^{\lambda}$: a larger $\lambda$ leads to larger regularization error, while reduces the approximation and generalization errors.
@@
A function implemented by a neural network $h:{\mathbb{R}}^{N_{0}}\rightarrow{\mathbb{R}}^{N_{L+1}}$ is defined by
@@
(4.3) ${\mathscr{H}}_{t}\coloneq\left\{h_{t}\in\mathrm{conv}(N(L,S)):\begin{aligned} &\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\leq\bar{B},\penalty 10000\ \inf_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\geq\underline{B},\\ &\max_{1\leq k\leq d}\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}|D_{k}h_{t}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B},\end{aligned}\right\}.$
where $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d}{d+4}})$. Let $\widehat{h}_{t}^{\lambda}$ be the gradient penalized empirical risk minimizer defined as (3.21), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequalities hold:
${\mathbb{E}}\Big[\|\nabla\log\widehat{h}_{t}^{\lambda}-\nabla\log h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\leq C\sigma_{T-t}^{-8}n^{-\frac{1}{d+4}},$
provided that the regularization parameter $\lambda$ is set as $\lambda={\mathcal{O}}(n^{-\frac{1}{d+4}})$, where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$.
The proof of Theorem 4.3 is deferred to Appendix C.4. This theorem demonstrates that the $L^{2}$-error of the Doob’s guidance estimator (3.22) converges to the exact Doob’s guidance in (3.14) as the sample size increases, provided that the size of the neural network is appropriately chosen. However, since the prefactor $\sigma_{T-t}^{-8}$ diverges as $t$ approaches the terminal time $T$, early stopping in controllable diffusion models (3.23) is required to ensure the validity of the Doob’s guidance estimator.

### v2

@@ 4.5  A summary of computing procedure
where $0\leq k\leq K-1$, the pre-trained reference score estimator $\widehat{\mathbf{s}}$ is defined as (2.5), and the Doob’s guidance estimator $\widehat{\mathbf{g}}^{\lambda}$ is defined as (4.7).
We apply post-processing to the generated particle $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$ to ensure numerical stability and facilitate the theoretical analysis presented in Theorem 5.6 (Lee et al., 2023a; Chen et al., 2023a). First, we assume the target distribution $q_{0}$ is concentrated on a domain centered at the origin, such as a distribution with compact support (Assumption 1) or with light tails. Consequently, we introduce a truncation operator to the particles obtained from the controllable diffusion model (4.8). Second, because the controllable diffusion process is terminated at an early-stopping time $T_{0}$, there exists a mean shift between the target distribution $q_{0}$ and the early-stopping distribution $q_{T_{0}}\approx\widehat{q}_{T-T_{0}}$, as indicated by (2.2). To mitigate this drift, we employ a scaling operator. Specifically, for $R>0$, we define a truncation operator ${\mathcal{T}}_{R}:\mathbf{z}\mapsto\mathbf{z}{\mathbbm{1}}_{B({\mathbf{0}},R)}(\mathbf{z})$ and a scaling operator ${\mathcal{M}}:\mathbf{z}\mapsto\mu_{T_{0}}^{-1}\mathbf{z}$. The final processed particle is defined as
@@
Input: Reference score estimator $\widehat{\mathbf{s}}$, the weight function $w$, the regularization parameter $\lambda$, the step size $h$, and the number of steps $K$.
@@
2 Estimate Doob’s $h$-function by $\widehat{h}_{t}^{\lambda}$ via gradient regularized Doob’s matching (4.6).
@@
6   Evaluate the reference score: $\widehat{\mathbf{s}}_{k}\leftarrow\widehat{\mathbf{s}}(kh,\widehat{\mathbf{Z}}_{kh}^{\leftarrow})$.
@@
Algorithm 1 Inference-time alignment via variationally stable Doob’s matching
@@ 5  Convergence Analysis
In this section, we derive a non-asymptotic convergence rate for the variationally stable Doob’s matching (4.6) and the induced controllable diffusion model (4.8). Furthermore, we demonstrate that this convergence rate mitigates the curse of dimensionality under mild assumptions.
@@ 5.1  Assumptions
The support of the target distribution $q_{0}$ is a compact set contained within the hypercube $\{\mathbf{x}_{0}\in{\mathbb{R}}^{d}:\|\mathbf{x}_{0}\|_{\infty}\leq 1\}$.
@@
Assumption 2 implies that the reference distribution $p_{0}$ and the tilted distribution $q_{0}$ satisfy mutual absolute continuity, ensuring that their supports coincide (i.e., $\mathrm{supp}(p_{0})=\mathrm{supp}(q_{0})$). This condition is crucial for establishing the regularity of Doob’s $h$-function. Furthermore, the ratio $\kappa\coloneq\bar{B}/\underline{B}$ serves as a condition number that characterizes the difficulty of the controllable diffusion task, as discussed in the context of posterior sampling by Purohit et al. (2025); Ding et al. (2024); Chang et al. (2025b).
Under Assumptions 1 and 2, we establish the regularity properties of Doob’s $h$-function defined in (3.10).
@@
The proof of Proposition 5.1 is deferred to Appendix C. It is worth noting that Proposition 5.1 relies solely on the boundedness of the weight function $w$, without requiring the existence or smoothness of its gradients. Nevertheless, we establish that Doob’s $h$-function admits bounded derivatives. This result stems from the definition of Doob’s $h$-function as a posterior expectation under a Gaussian likelihood; the inherent smoothness of the Gaussian kernel endows the posterior expectation with strong regularity properties.
Assumption 3 (Reference score estimation error).
The reference score estimator $\widehat{\mathbf{s}}$ defined in (2.5) satisfies the following error bound:
$\frac{1}{T}\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{P}}}\left[\|\widehat{\mathbf{s}}(kh,\mathbf{X}_{kh}^{\leftarrow})-\nabla\log p_{T-kh}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right]\leq\varepsilon_{\mathrm{ref}}^{2}.$
Assumption 3 requires the $L^{2}$-error of the reference score estimator $\widehat{\mathbf{s}}$ to be bounded with respect to the reference path measure $\mathbb{P}$}. In our setting, where numerous samples from the reference distribution $p_{0}$ are available, estimators satisfying this bound can be obtained via implicit score matching (Hyvärinen, 2005), sliced score matching (Song et al., 2020), or denoising score matching (Vincent, 2011). While one can derive explicit bounds of reference score matching as Tang and Yang (2024); Oko et al. (2023); Fu et al. (2024); Ding et al. (2025a); Yakovlev and Puchkin (2025a); Yakovlev et al. (2025) using non-parametric regression theory for deep neural networks (Bauer and Kohler, 2019; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Jiao et al., 2023), we adopt this condition to maintain clarity of presentation, following the convention of Lee et al. (2023a); Chen et al. (2023a); Beyler and Bach (2025); Kremling et al. (2025).
@@ 5.2  Error bounds for the Doob’s guidance estimator
We begin by introducing the concept of Vapnik-Chervonenkis (VC) dimension (Vapnik and Chervonenkis, 1971; Anthony et al., 1999; Bartlett et al., 2019), which measures the complexity of a function class.
Definition 1 (VC-dimension).
Let $\mathscr{H}$} be a class of functions mapping from $\mathcal{X}$} to $\mathbb{R}$}. For any num-negative integer $m$, the growth function of $\mathscr{H}$} is defined as
$\Pi_{{\mathscr{H}}}(m)\coloneq\max_{x_{1},\ldots,x_{m}\in{\mathcal{X}}}|\{(\mathrm{sgn}\,h(x_{1}),\ldots,\mathrm{sgn}\,h(x_{m})):h\in{\mathscr{H}}\}|.$
We say $\mathscr{H}$} shatters the set $\{x_{1},\ldots,x_{m}\}$, if
$|\{(\mathrm{sgn}\,h(x_{1}),\ldots,\mathrm{sgn}\,h(x_{m})):h\in{\mathscr{H}}\}|=2^{m}.$
The Vapnik-Chervonenkis dimension of $\mathscr{H}$}, denoted by $\mathrm{VCdim}({\mathscr{H}})$, is the size of the largest shattered set, i.e., the largest $m$ such that $\Pi_{{\mathscr{H}}}(m)=2^{m}$.
To simplify notation, we define the gradient classes and their associated VC-dimensions. For a differentiable hypothesis class $\mathscr{H}$} consisting of functions mapping from ${\mathbb{R}}^{d}$ to $\mathbb{R}$}, the VC-dimension of the gradient hypothesis class is defined as
$\mathrm{VCdim}(\nabla{\mathscr{H}})\coloneq\max_{1\leq k\leq d}\mathrm{VCdim}(D_{k}{\mathscr{H}}),\quad D_{k}{\mathscr{H}}\coloneq\{D_{k}h:h\in{\mathscr{H}}\},$
where $D_{k}$ represents the derivative with respect to the $k$-th entry of the input.
The following lemma provides an oracle inequality for the variationally stable Doob’s matching (4.6).
@@
Suppose Assumptions 1 and 2 hold. Let $t\in(0,T)$ and let ${\mathscr{H}}_{t}$ be a hypothesis class. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequalities hold:
${\mathbb{E}}\Big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]$ $\lesssim\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big\}}_{\text{(I)}}$
$\quad+\underbrace{\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}}_{\text{(II)}}+\underbrace{\frac{\lambda^{2}d\bar{B}^{2}}{\sigma_{T-t}^{8}}}_{\text{(III)}},$
${\mathbb{E}}\Big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]$ $\lesssim\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{\frac{1}{\lambda}\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big\}}_{\text{(I))}}$
$\quad+\underbrace{\frac{\bar{B}^{2}}{\lambda}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}}_{\text{(II)}}+\underbrace{\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{8}}}_{\text{(III)}},$
@@
The proof of Lemma 5.2 is deferred to Appendix D. Both oracle inequalities for Doob’s $h$-function and its gradient decompose the error into three components: approximation error, generalization error, and regularization gap.
@@
    The approximation error is defined as the minimal $H^{1}$-distance between functions in the hypothesis class ${\mathscr{H}}_{t}$ and the ground-truth Doob’s $h$-function $h_{t}^{*}$, measuring the approximation capability of ${\mathscr{H}}_{t}$.
@@
    The regularization gap is introduced by the gradient regularization in the objective functional, which causes the minimizer of the variationally stable objective (4.3) to deviate from the ground-truth Doob’s $h$-function $h_{t}^{*}$ (3.10). This gap has been analyzed in Proposition 4.2.
@@
Lemma 5.2 is analogous to the oracle inequality found in regression problems. Let $\widehat{h}_{t}$ be the vanilla estimator estimated by minimizing (4.5) over the hypothesis class ${\mathscr{H}}_{t}$. Informally, the following oracle inequality holds:
(5.1) ${\mathbb{E}}\Big[\|\widehat{h}_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\lesssim\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}}_{\text{approximation}}+\underbrace{\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}}_{\text{generalization}}.$
@@
    In vanilla regression, the generalization error in (5.1) depends only on the complexity of the hypothesis class. In contrast, the generalization error bounds in Lemma 5.2 also depend on the complexity of the derivative classes $\nabla{\mathscr{H}}_{t}$. This occurs because the objective functional of the gradient-regularized Doob’s matching (4.3) includes the gradient norm term. Consequently, the error from finite-sample approximation is influenced not only by the complexity of the hypothesis class but also by that of the derivative class.
@@
    The most significant difference lies in the regularization error. If we focus solely on the oracle inequality for $\widehat{h}_{t}^{\lambda}$, letting $\lambda$ go to zero reduces the expression to the vanilla regression oracle inequality (5.1). However, the bound for the gradient $\nabla\widehat{h}_{t}^{\lambda}$ diverges as the regularization parameter $\lambda$ approaches zero. This highlights the key advantage of our gradient-regularized method: the gradient-regularized is essential for guaranteeing simultaneous convergence of both the estimator value and its gradient. Additionally, there exists a trade-off with respect to $\lambda$ in the oracle inequality for $\widehat{h}_{t}^{\lambda}$: a larger $\lambda$ leads to larger regularization error, while reduces the approximation and generalization errors.
@@
A function implemented by a neural network $h:{\mathbb{R}}^{N_{0}}\to{\mathbb{R}}^{N_{L+1}}$ is defined by
@@
(5.2) ${\mathscr{H}}_{t}\coloneq\left\{h_{t}\in N(L,S):\begin{aligned} &\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\leq\bar{B},\penalty 10000\ \inf_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\geq\underline{B},\\ &\max_{1\leq k\leq d}\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}|D_{k}h_{t}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B}\end{aligned}\right\},$
where $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d}{d+8}})$. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequality holds:
${\mathbb{E}}\Big[\|\nabla\log\widehat{h}_{t}^{\lambda}-\nabla\log h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\leq C\sigma_{T-t}^{-8}n^{-\frac{2}{d+8}}\log^{4}n,$
provided that the regularization parameter $\lambda$ is set as $\lambda={\mathcal{O}}(n^{-\frac{2}{d+8}})$, where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$.
The proof of Theorem 5.3 is deferred to Appendix D. This theorem demonstrates that the $L^{2}$-error of the Doob’s guidance estimator (4.7) converges to the exact Doob’s guidance in (3.14) as the sample size increases, provided that the size of the neural network is appropriately chosen. However, since the prefactor $\sigma_{T-t}^{-8}$ diverges as $t$ approaches the terminal time $T$, early stopping in controllable diffusion models (4.8) is required to ensure the validity of the Doob’s guidance estimator.
Remark 5.4 (Comparisons with previous work).
Simultaneous estimation of a function and its gradient using deep neural network has been investigated by Ding et al. (2025b). The most important distinction in our work lies in the elimination of the convexity assumption on the hypothesis class. Specifically, Ding et al. (2025b, Lemma 7) propose an oracle inequality under the assumption that the hypothesis class is convex. Furthermore, Ding et al. (2025b, Theorem 3) requires the estimator to be a minimizer of the gradient-regularized empirical risk over the convex hull of a neural network class, which is intractable in practice. In contrast, Lemma 5.2 eliminates the requirement of convexity for the hypothesis class, and Theorem 5.3 removes the need for the convex hull of the neural network class. This aligns the theoretical analysis more closely with practical computing.